answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What is the probability of selecting a red even card in a deck of 52?

The probability of selecting a red card is 26 in 52 or 1 in 2. The probability of selecting an even card is 20 in 52 or 5 in 13. The probability, therefore, of selecting a red even card is 1 in 2 times 5 in 13 or 5 in 26.

What is the formula of experimental probability?

If an event occurs in n trials out of N experiments than the experimental probability of that event is n/N.

What is the probability of throwing the dice and getting a total of 5 on the first throw and getting a total of 10 on the second throw?

You have not given any information about the shape and numbering of the dice nor how many are thrown each time. Without that information it is not possible to answer the question.

How do you do the probability of spinning a 3?

In general, you cannot. The answer depends on the shape of the spinner - how many sides it has and what numbers are on it.

What is the probability that at least one of the dice will roll three if you roll four dice at the same time?

The probability that one die will not be a 3 is 5/6

The probability that all will not be 3 is ((5/6)4)

If you roll 4 dice, there are 5 possible outcomes (for this problem).

There will be NO 3's, there will be one 3, there will be two 3's. there will be three 3's, or, there will be four 3's. Add all possible probabilities = 1. At least one meanse either 1,2,3, or 4 3's but NOT one.

So the prob of at least one = total prob - prob of NO

1- (5/6)4 = 1 - 625/1296

= 671/1296 <-- answer

How many possible ways to answer three true or false questions?

2 possibilities: true or false

3 questions

2^3

2*2*2 = 8 ways

Is continuous distribution convert in discrete distribution?

If the question is asking if a continuous distribution can be converted to a discrete distribution, the answer is yes.

Your age has a continuous distribution but in most cases, the information is recorded and analysed as if it were the whole number of years - a discrete distribution.

How do you create five scores with a mean of 10 and a standard deviation of 0?

Since the standard deviation is zero, the scores are all the same. And, since their mean is 10, they must all be 10.

How useful probability in statistics?

Statistics is based on the theoretical foundation of probability.

What is the probability of rolling a 6 on two six sided dice?

The probability of rolling a six on either (or both) die is 11/36.

What is the probability of drawing a diamond and a black seven in a deck of 52 cards?

# of diamond card = 13

# of black 7 card = 2

P(diamond and black 7) = 13/52 * 2/51 = 1/4 * 2/51

Answer = 1/102

What is relations between representative sample and population?

A representative sample is a randomly selected subset of the population.

What is the probability of getting tails 4 times in a row when you flip a coin 4 times?

The probability of getting tails in any one flip is 1/2, so 1/2 x 1/2 x 1/2 x 1/2 = 1/16, or 6.25%

How many different ways can you arrange the letters of the word party?

120 ways, 5! (5 factorial)

for first place you have 5 candidates,

for second place you have 4 candidates,

for third place you have 3 candidates,

for fourth place you have 2 candidates,

for last place you have 1 candidates,

5 x 4 x 3 x 2 x 1 = 120 different ways

What is the probability that three F2 seed chosen at random will include at least one yellow seed?

The probability that three F2 seeds chosen from Mendel's study group will have at least one yellow seed is 63/64. It would be very rare to get three green seeds.

How many possible outcomes in a choice of muffin or toast with coffee milk or juice?

just multiply them together 2*3=6

muffin-(coffee) OR (milk) OR (juice): three possibilities

toast- (coffee) OR (milk) OR (juice): three possibilities

6 total